OpenStudy (ddcamp):

Could someone review this proof for me? Prove: A^n - B^n = (A-B)×(A+B)^n-1 (A+B)^n-1 = (A^n-1)+(A^n-2)(B^1)+(A^n-3)(B^2)+...+(A^1)(B^n-2)+(B^n-1) (A-B)(A+B)^n-1 = [(A^n)+(A^n-1)(B^1)+(A^n-2)(B^2)+...+(A^2)(B^n-2)+(A^1)(B^n-1)] - [(A^n-1)(B^1)+(A^n-2)(B^2)+(A^n-3)(B^3)+...+(A^1)(B^n-1)+(B^n)] Cancel out like terms: (A-B)(A+B)^n-1= (A^n)-(B^n)

OpenStudy (ddcamp):

Prove: \[A^n - B^n = (A-B)(A+B)^{ n-1}\] Binomial exponents: \[(A+B)^{n-1} = (A^{n-1})+(A^{n-2})(B^1)+(A^{n-3})(B^2)+...+(A^1)(B^{n-2})+(B^{n-1})\] Factoring in (A-B): \[(A-B)(A+B)^{n-1} = [(A^n)+(A^{n-1})(B^1)+(A^{n-2})(B^2)+...+(A^2)(B^{n-2})+(A^1)(B^{n-1})]\] \[- [(A^{n-1})(B^1)+(A^{n-2})(B^2)+(A^{n-3})(B^3)+...+(A^1)(B^{n-1})+(B^n)]\] Cancelling like terms: \[(A-B)(A+B)^{n-1}= A^n-B^n\]

OpenStudy (ddcamp):

I think it's right but I would like someone else to make sure. I'd ask one of my proffs but I won't be seeing any of them for another 2 months.

OpenStudy (anonymous):

Do you know how to prove things using induction?

OpenStudy (ddcamp):

Is that the "Prove it for n=1" then "Assume it's true for n and prove it for n+1"?

OpenStudy (anonymous):

Yup, that's it.

OpenStudy (anonymous):

I don't know what kind of rigor is required here, but that's how I'd go about proving this.

OpenStudy (anonymous):

the (n+1) case will yield a part of the original base case. My assumption is that when going through that step you'll wide up with a bunch of stuff that'll cancel out and arrive at the conclusion that they are indeed equal.

OpenStudy (ddcamp):

Cool, thanks!

OpenStudy (anonymous):

No problem! Induction is really such an awesome tool. :D